OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The average of two test scores is 79. If one test score is eleven less than the other, what are the two test scores?

OpenStudy (anonymous):

\[(x+(x-11))\div2=79\]

OpenStudy (anonymous):

2x-11=158 2x=169 x=84.5? x-11=73.5? h/o 84.5+73.5=158 158/2=79. hmmm...weird test scores

OpenStudy (anonymous):

Very weird... thanks for your help! :)